A circular cone piece of cardboard with a diameter of 1 m will be made into a conical hat 44 cm high by cutting a sector off and joining the edges to form a cone. Determine the anglein degree subtended by the sector removed.


1
Expert's answer
2021-05-25T10:19:14-0400
"r=100\ cm/2=50\ cm, h=44\ cm"


From the right triangle


"x=\sqrt{r^2-h^2}=\sqrt{50^2-44^2}=2\sqrt{141}(cm)"

Let:

"C_1=" circumference of the circle


"C_1=2\pi r"

"C_2=" circumference of the base of the cone


"C_2=2\pi x"

"C=" length of the arc


"C=r\theta"



Then


"C=C_1-C_2"

"r\theta=2\pi r-2\pi x"

"\theta=\dfrac{2\pi(r-x)}{r}"

"\theta=2\pi(1-\dfrac{x}{r})"

"\theta\degree=2\pi(1-\dfrac{x}{r})(\dfrac{360\degree}{2\pi})"

"\theta\degree=(1-\dfrac{x}{r})360\degree"

"\theta\degree=(1-\dfrac{2\sqrt{141}}{50})360\degree\approx189\degree"